tan36+tan9+tan36tan9=1 prow that
  please   ans the sum
 

Grade:11

1 Answers

ADARSH SANKAR R S
7 years ago
As tan 45 = 1 = tan ( 36 + 9 ) = 1
= ( tan 36 + tan 9 ) / ( 1 - tan 9 tan 36 ) = 1
= tan 36 + tan 9 = 1 - tan 9 tan 36
= tan 36 + tan 6 + tan 9 tan 36 = 1
Hence Proved......